OpenStudy (anonymous):

integral (3x^2 +x+4)/(x^3+x)

OpenStudy (anonymous):

\[\int\limits (3x^2 +x+4)/(x^3+x)\]

OpenStudy (anonymous):

(3x^@+x+4)/[x(x^2+1)]=a/x+bx/(x^2+1)

OpenStudy (anonymous):

find a,b

OpenStudy (anonymous):

well, I dont know when to use Bx. can you please explain it?

OpenStudy (anonymous):

i used Bx + C and got a result that checks out i got 4 / x^2 + 1 /(x^2 + 1) - x /( x^2 + 1)

OpenStudy (anonymous):

when ur denominator is a quadratic then use bx+c